# Intelligent Surfaces Empowered Wireless Network: Recent Advances and The Road to 6G

### 15 Jan 2024 v2

@@ II-A3 Beam Training
In Table III, we summarize the main approaches for the IS beam training designs under both the far-/near-field channel models. Despite these existing works, there are several practical challenges that need to be solved in IS beam training. For example, it is practically important to analyze the impact of hardware constraints/imperfections (e.g., discrete phase shift, amplitude-phase dependent IS reflection) on the IS beam training performance and develop robust beam training schemes. Besides, accurate synchronization between the BS and IS is also important to achieve joint beam training at the BS and IS. To fully exploit the potentials of IS in 6G systems, future research may also consider the joint optimization of IS deployment and beamforming, coordinated beamforming of multiple ISs, statistical channel-based IS reflection optimization for coverage enhancement, as well as machine learning-aided IS channel recovery and reflection optimization.
@@ II-B Deployment
Figure 9: An illustration of IS-integrated BS, where the IS elements and BS antennas are integrated in the same antenna radome.
@@ II-B1 BS-Side IS
The existing literature has mainly considered deploying ISs near user terminals to improve their performance. However, this approach may incur a high cost since ISs need to be densely deployed in the environment to cater to random user locations. Moreover, even with dense IS deployment, each user can only enjoy the passive beamforming gain brought by its nearby IS (as its signals reflected by other far-away ISs are generally much weaker due to more significant path loss). In [94], the authors have characterized the capacity regions achievable by two IS deployment strategies with the IS/ISs deployed near the BS and each of distributed users, respectively, and showed the superiority of the former over the latter under the same total number of IS REs.
@@ II-B2 Multi-IS Deployment
Despite the above appealing advantages of the BS-side IS, a blockage-free BS-user link via any BS-side IS may not be available in practice, since the reflection paths by BS-side ISs only may fail to bypass the dense and scattered obstacles in complex environments (e.g., an indoor environment shown in Fig. 10(a)). To overcome this limitation, recent works[48, 98, 99, 100, 101, 102] have proposed a multi-IS-reflection aided wireless system, where multiple distributed ISs are deployed in the environment to create more available cascaded LoS paths between the BS and remote users, thereby significantly enhancing the wireless coverage, as shown in Fig. 10(b). To achieve this purpose, how to properly deploy multiple ISs to cater to the practical environment is a critical problem of high importance.
@@
However, the communication performance in [104] was only evaluated in terms of the number of signal reflections per multi-IS link, which may not accurately indicate the actual performance of each user. To tackle this issue, the authors in [104] proposed an enhanced graph-based system modeling approach which is able to characterize the SNR performance at any user location with a given multi-IS deployment solution. Moreover, to more effectively overcome the multiplicative path loss over the multi-reflection LoS links, the joint use of active and passive ISs was studied in [104]. The results in [104] revealed that incorporating active ISs is beneficial to reduce the total deployment cost for achieving a given SNR target in the target region by reducing the number of passive ISs required.
@@ II-B3 Relay-side ISs
When deployed in wireless networks, passive IS and active relay differ in many aspects such as energy/spectral efficiency, hardware cost, processing complexity, and coverage range [106, 107]. Specifically, passive IS can operate in full duplex and enhance the communication performance within its local coverage cost-effectively; while active relay typically operates in half duplex and is able to achieve broader coverage at the expense of more power consumption [68, 108, 109]. To harness the advantages of both, the relay-side IS deployment strategy has been considered. However, most of the existing works on relay-side IS simply considered adding ISs near the active relay [110, 111, 112, 113, 114], which needs to deploy ISs and relays separately and also enable their cooperation. This thus incurs additional deployment cost and signaling overhead. In contrast, an alternative relay-side IS deployment scheme was proposed in [105] to integrate the active relay directly into the IS, by exploiting the new role of the IS controller as an active relay, as shown in Fig. 11. Specifically, IS-integrated relay significantly enhances coverage performance compared to conventional ISs (with passive signal reflection/refraction only), by enabling the IS controller to actively relay information without additional deployment or hardware cost. Moreover, similar to IS-integrated BS, IS-integrated relay also has a considerably shorter distance between IS REs and the active relay (see Fig. 11). This yields several advantages such as reduced product-distance path loss (equivalently, increased effective channel gain), improved path diversity, expanded coverage range, and reduced signaling/information exchange overhead.
@@
However, the channel estimation for the IS-integrated relay-aided communication system becomes more complicated than that for the conventional IS-aided or relay-aided communication systems due to the increased number of channel coefficients and the near-field channel response between IS elements and relay antennas. To address this issue, a practical channel estimation scheme was proposed in [105], where the cascaded/direct CSI of the BS-relay link and the relay-user link is estimated in parallel at the BS and user, respectively, based on the pilot signals sent from the relay. For practicality and to reduce training overhead, a codebook-based reflection design tailored for the IS-integrated relay may be considered, which needs to account for the mixed near- and far-field effect between relay/IS and BS/users (see Fig. 11). On the other hand, half-duplex relay is considered in [105] for ease of implementation, where two-phase relaying transmission is required and the IS can enhance both BS-relay and relay-user links by exploiting its time selectivity property over two phases. To boost transmission efficiency, the full-duplex relay could be considered for one-phase relaying transmission. However, this may introduce additional self-interference, necessitating sophisticated signal processing techniques to properly manage the interference, and the IS reflection would need to balance the transmission performance between BS-relay and relay-user links.
@@ II-B4 Aerial ISs
In addition to being deployed terrestrially, ISs can also be deployed on aerial platforms such as UAVs or balloons to achieve signal enhancement by providing additional LoS air-to-ground or air-to-air links, especially for post-disaster or scarcely covered areas [115, 116]. To fulfill this application, the joint optimization of the aerial IS placement and the 3D passive beamforming was considered in [117, 118]. Despite exploiting the additional design degree-of-freedom (DoF) provided by the vertical dimension, the enhanced performance resulting from aerial ISs is mainly constrained by the limited onboard energy and computational capabilities. To overcome these challenges, practical energy-efficient resource allocation and placement/trajectory optimization schemes have been proposed for designing aerial ISs [119, 120], where DL-based algorithms exhibit potentials for solving non-convex optimization problems with coupled variables more efficiently as compared to conventional optimization algorithms [121, 122]. However, the signal transmitted on aerial links could suffer from high path loss, especially for high frequency signals under long-distance transmission. To address this issue, aerial relaying architectures were studied in [123, 124] where UAV-mounted ISs relay the transmitted signals from aerial platforms at higher altitudes. Nevertheless, to be more widely used in practice, further investigation on new architectures and protocols are needed to address the challenge of accurate and timely CSI acquisition arising from the mobility of aerial ISs.
@@ II-B5 Roadside and Vehicle-side ISs
TABLE IV: Comparison of Roadside IS and Vehicle-side IS [129]
@@
An alternative strategy is to install a vehicle-side IS at each vehicle to constantly improve communication performance between the remote BS and the IS-aided user [130]. Different from roadside IS deployment, vehicle-side IS deployment only requires one IS per vehicle. In this case, the vehicle-side IS can easily associate with the onboard users constantly and much less frequent IS-BS handover is triggered. The comparisons between roadside and vehicle-side IS deployment are summarized in Table IV. Note that both roadside and vehicle-side ISs can be based on either signal reflection or refraction, depending on the deployment conditions. Moreover, to achieve high passive beamforming gain for both roadside and vehicle-side ISs, accurate CSI acquisition is essential yet practically challenging. This is due to the high-dimensional CSI associated with a large number of IS elements and the short channel coherence time due to high mobility. As such, instead of full channel estimation/acquisition, it is practically appealing to perform beam alignment to efficiently track the fast time-varying cascaded user-IS-BS channel with low training overhead, which deserves further investigation.
@@ II-C Modulation
From a network functionality and information transmission perspective, we can classify ISs into two types depending on their roles in the network: relay-type IS and transmitter-type IS [131]. Specifically, an individual IS can function as either a relay or a transmitter, depending on its functionality and the specific use cases. Relay-type IS, as depicted in Fig. 12 and discussed in the previous subsections, is strategically positioned between the source (e.g., BS) and the destination (e.g., users) to enhance communication performance between them, especially when the direct link is considerably weak or obstructed. On the other hand, transmitter-type IS is an integral part of the transmitter, participating in its signaling and modulation processes, as illustrated in Fig. 12. In this subsection, we focus on the transmitter-type IS and provide an overview of two types of transmitter-type IS-based modulation: amplitude/phase modulation (APM) and index modulation (IM).
@@ II-C2 IS for Index Modulation
Despite these promising recent advances in IS-based modulation designs, significant challenges remain to fully unlock their potential for future wireless networks. Using transmitter-type ISs is beneficial in designing transmitters with simplified hardware. However, in this case, the resulted transmitter may become bulky, considering the integration of the IS into the structure. Furthermore, generating QAM signals is not straightforward, and the signal coverage is usually limited. As a result, IS-based transmitter design is an active area of research, particularly for effective practical implementation. Additionally, the low data rate of reflection modulation designs is a significant issue for their effective data transmission. In this regard, designing novel IM-based IS systems that can reliably embed more information bits into IS is an important direction for future work.
@@ II-D Sensing and ISAC
ISAC has been regarded as a promising technique for the 6G wireless network, where the sensing function and the communication function can be carried out simultaneously by sharing the same spectrum and hardware facilities [151, 152]. However, ISAC systems still face practical challenges. For example, sensing coverage is significantly smaller than communication coverage due to severe path loss in the BS-IS-target-IS-BS cascaded echo link. Although the sensing coverage can be guaranteed by deploying more BSs, this method incurs high hardware cost and energy consumption. On the other hand, hardware sharing of communication and sensing transceivers increases the design and implementation complexity, while spectrum sharing of communication and sensing signals results in interference and performance trade-off. Fortunately, the use of IS can achieve higher spectral/energy efficiency for wireless communication[153, 154], as well as higher accuracy for wireless sensing, thus promising to enable ISAC for 6G [155, 156]. Specifically, the IS can establish LoS channels to enhance both sensing/communication coverage, provide flexible passive beamforming to balance their performance trade-off, and reduce hardware cost and energy consumption compared to dense deployment of BSs.

### 24 Mar 2024 v3

@@ II-A3 Beam Training
In Table IV, we summarize the main approaches for the IS beam training designs under both the far-/near-field channel models. Despite these existing works, there are several practical challenges that need to be tackled in IS beam training. For example, it is practically important to analyze the impact of hardware constraints/imperfections (e.g., discrete phase shift, amplitude-phase dependent IS reflection) on the IS beam training performance and develop robust beam training schemes. Besides, accurate synchronization between the BS and IS is also important to achieve joint beam training at the BS and IS. To fully exploit the potentials of IS in 6G systems, future research may also consider the joint optimization of IS deployment and beamforming, coordinated beamforming of multiple ISs, statistical channel-based IS reflection optimization for coverage enhancement, as well as machine learning-aided IS channel recovery and reflection optimization.
Summary: In this subsection, we have provided an overview of the state-of-the-art approaches for IS reflection/beamforming design based on received signal power measurements. For sub-6 GHz bands, the blind reflection/beamforming approach and channel recovery-based approach are both applicable. The blind reflection/beamforming approach does not require explicit CSI and has a low computational complexity. However, to achieve a satisfactory performance gain, this approach usually needs a sufficiently large number of power measurements for extracting the useful statistics in CSI, which inevitably incurs high training overhead. In contrast, the channel recovery-based approach requires less power measurements but higher computational complexity for channel estimation and IS reflection design. Both approaches can be applied to real-time IS reflection design for improving the instantaneous channels between the BS and users, under the condition of slow-fading channels. While for fast-fading channels, the two approaches are still feasible because they can exploit the statistical CSI or the deterministic part of time-varying wireless channels. As such, the corresponding IS reflection design can improve the spatial power distribution of channels between the BS and users over a long period. For high-frequency bands (e.g., mmWave), beam training is a more efficient approach for designing IS beamforming by effectively exploiting the channel paths sparsity in the spatial domain. Since this approach does not require the explicit CSI of IS-involved links and only selects the best directional beam from a designed codebook, its required training overhead and computational complexity are much lower than those of the blind reflection/beamforming approach and channel recovery-based approach. In addition to the pencil-like beamforming based on instantaneous channels, the wide beam of IS can also be designed based on the statistical channel knowledge/deterministic channel components to improve the long-term coverage performance over a target region.
@@ II-B Deployment
Figure 9: An illustration of IS-integrated BS, where the IS elements and BS antennas are integrated within the same antenna radome.
@@ II-B1 BS-Side IS
The existing literature has mainly considered deploying ISs near user terminals to improve their performance. However, this approach may incur a high cost since ISs need to be densely deployed in the environment to cater to random user locations. Moreover, even with dense IS deployment, each user can only enjoy the passive beamforming gain brought by its nearby IS (as its signals reflected by other far-away ISs are generally much weaker due to more significant path loss). In [162], the authors have characterized the capacity regions achievable by two IS deployment strategies with the IS/ISs deployed near the BS and each of the distributed users, respectively, and showed the superiority of the former over the latter under the same total number of IS REs.
@@
Note that a BS-side IS is similar to a reflectarray as they both reflect signals from active antennas to attain the passive beamforming gain. Nonetheless, there are also differences between them. On one hand, compared to the reflectarray, the deployment of ISs is more flexible, which can be placed in the vicinity of the BS or integrated into the antenna radome of the BS. In contrast, a conventional reflectarray is usually placed at the backside of active antenna. On the other hand, reflectarrays are usually composed of a single planar surface mounted with reflecting elements. However, BS-side ISs can deploy multiple reflective surfaces with different positions and orientations around the BS [162, 163], which can exploit their single and double reflection gains to further improve communication performance. As such, the conventional reflectarray can be regarded as a special case of the BS-side IS by deploying a single reflective surface at the backside of active antenna only.
@@ II-B2 Multi-IS Deployment
Despite the above appealing advantages of the BS-side IS, a blockage-free BS-user link via any BS-side IS may not be available in practice, since the reflection paths by BS-side ISs only may fail to bypass the dense and scattered obstacles in complex environments (e.g., an indoor environment shown in Fig. 10(a)). To overcome this limitation, recent works[103, 166, 167, 168, 169, 170, 171, 127] have proposed a multi-IS-reflection aided wireless system, where multiple distributed ISs are deployed in the environment to create more available cascaded LoS paths between the BS and remote users, thereby significantly enhancing the wireless coverage, as shown in Fig. 10(b). To achieve this purpose, how to properly deploy multiple ISs to cater to the practical environment is a critical problem of high importance.
@@
However, the communication performance in [173] was only evaluated in terms of the number of signal reflections per multi-IS link, which may not accurately indicate the actual performance of each user. To tackle this issue, the authors in [173] proposed an enhanced graph-based system modeling approach which is able to characterize the SNR performance at any user location with a given multi-IS deployment solution. Moreover, to more effectively overcome the multiplicative path loss over the multi-reflection LoS links, the joint use of active and passive ISs was studied in [173]. The results in [173] revealed that incorporating active ISs is beneficial to reduce the total deployment cost for achieving a given SNR target in the target region by dramatically reducing the number of passive ISs required.
@@ II-B3 Relay-side ISs
When deployed in wireless networks, passive IS and active relay differ in many aspects such as energy/spectral efficiency, hardware cost, processing complexity, and coverage range [175, 176]. Specifically, passive IS can operate in full duplex and enhance the communication performance within its local coverage cost-effectively; while active relay typically operates in half duplex and is able to achieve broader coverage at the expense of more power consumption [115, 177, 178]. To harness their combined advantages, the relay-side IS deployment strategy has been proposed in some prior works. However, most of the existing works on relay-side IS simply add ISs to the relay system and deploy them separately [179, 180, 181, 182, 183], which thus incurs additional deployment cost and signaling overhead. In contrast, an alternative relay-side IS deployment scheme was proposed in [174] to integrate the active relay directly into the IS, by exploiting the new role of the IS controller as an active relay, as shown in Fig. 11. Specifically, IS-integrated relay significantly enhances coverage performance compared to conventional ISs (with passive signal reflection/refraction only), by enabling the IS controller to actively relay information without additional deployment or hardware cost. Moreover, similar to IS-integrated BS, IS-integrated relay also has a considerably shorter distance between IS REs and the active relay (see Fig. 11), which yields several advantages such as reduced product-distance path loss (equivalently, increased effective channel gain), improved path diversity, expanded coverage range, and reduced signaling/information exchange overhead.
@@
However, the channel estimation for the IS-integrated relay-aided communication system becomes more challenging than that for the conventional IS-aided or relay-aided communication systems due to the increased number of channel coefficients and the near-field channel response between IS elements and relay antennas. To address this issue, a practical channel estimation scheme was proposed in [174], where the cascaded/direct CSI of the BS-relay link and the relay-user link is estimated in parallel at the BS and user, respectively, based on the pilot signals sent from the relay. For practicality and to reduce training overhead, a codebook-based reflection design tailored for the IS-integrated relay may be considered, which needs to account for the mixed near- and far-field effect between relay/IS and BS/users (see Fig. 11). On the other hand, half-duplex relay is considered in [174] for ease of implementation, where two-phase relaying transmission is required and the IS can enhance both BS-relay and relay-user links by exploiting its time selectivity property over two phases. To boost transmission efficiency, the full-duplex relay could be considered for one-phase relaying transmission. However, this may introduce additional self-interference, necessitating sophisticated signal processing techniques to properly manage the interference, and the IS reflection would need to balance the transmission performance between BS-relay and relay-user links.
@@ II-B4 Aerial ISs
In addition to being deployed terrestrially, ISs can also be deployed or mounted on aerial platforms such as UAVs or balloons to achieve signal enhancement by providing additional LoS air-to-ground or air-to-air links, especially for post-disaster or scarcely covered areas [36, 184, 185, 37, 186]. To fulfill this application, the joint optimization of the aerial IS placement and the 3D passive beamforming was considered in [187, 188]. Despite exploiting the additional design degree-of-freedom (DoF) provided by the vertical dimension, the enhanced performance resulting from aerial ISs is mainly constrained by the limited onboard energy and computational capabilities. To overcome these challenges, practical energy-efficient resource allocation and placement/trajectory optimization schemes have been proposed for designing aerial ISs [189], where DL-based algorithms exhibit potentials for solving non-convex optimization problems with coupled variables more efficiently as compared to conventional optimization algorithms [190]. However, the signal transmitted on aerial links could suffer from high path loss, especially for high frequency signals under long-distance transmission. To address this issue, aerial relaying architectures were studied in [191] where UAV-mounted ISs relay the transmitted signals from aerial platforms at higher altitudes.
Furthermore, the application of IS to other types of airborne systems such as low earth orbit (LEO) satellite communication is an emerging direction of high interest. Specifically, LEO satellite communication systems suffer from the large Doppler effect due to the high mobility of satellites, while its unfavorable effects on signal quality can be effectively mitigated by ISs [192, 38]. In addition, the qualities of satellite-terrestrial and inter-satellite links can be improved by utilizing ISs. For example, a new IS-aided satellite communication system with two-sided cooperative ISs has been proposed, where the communications between the LEO satellite and various ground nodes are enhanced by distributed ISs deployed near them [193]. The authors of [193] also proposed an efficient cooperative beamforming design and a practical transmission protocol to achieve distributed channel estimation and beam tracking, and demonstrated the substantial performance gain and great potential of the IS-aided satellite communication. Nevertheless, to be more widely used in practice, the effects of aerial IS’s mobility (e.g., trajectory drifts due to hardware imperfections) on system performance deserve further investigation.
@@ II-B5 Roadside and Vehicle-side ISs
TABLE V: Comparison of Roadside IS and Vehicle-side IS [198]
@@
An alternative strategy is to install a vehicle-side IS at each vehicle to constantly improve communication performance between the remote BS and the IS-aided user [199]. Different from roadside IS deployment, vehicle-side IS deployment only requires one IS per vehicle. In this case, the vehicle-side IS can easily associate with the onboard users constantly and much less frequent IS-BS handover is triggered. The comparisons between roadside and vehicle-side IS deployment are summarized in Table V. Note that both roadside and vehicle-side ISs can be based on either signal reflection or refraction, depending on the deployment conditions. Moreover, to achieve high passive beamforming gain for both roadside and vehicle-side ISs, accurate CSI acquisition is essential yet practically challenging. This is due to the high-dimensional CSI associated with a large number of IS elements and the short channel coherence time due to high mobility. As such, instead of full channel estimation/acquisition, it is practically appealing to perform beam alignment to efficiently track the fast time-varying cascaded user-IS-BS channel with low training overhead, which deserves further investigation.
Summary: In this subsection, we have reviewed the state-of-the-art IS deployment strategies, including BS-side ISs, multiple cooperative ISs, relay-side ISs, aerial ISs, roadside ISs and vehicle-side ISs, which are important for practical use and adaption to various wireless environments and application scenarios. From the perspective of network architecture, BS-side ISs can save the high cost of densely deploying user-side ISs and also provide better overall system performance. Apart from deploying ISs near BSs, the IS-integrated BS architecture can further decrease the product-distance path loss by integrating IS REs into the BS’s antenna radome, while the increased channel estimation complexity can be addressed by codebook-based IS reflection design. In complicated wireless environment with dense obstacles and scatterers, multiple ISs can be deployed to establish cascaded virtual LoS links by utilizing graph-based system modeling and optimization methods to improve the communication performance. For further performance enhancement, ISs can be deployed close to or integrated into active relays to combine the advantages of both. Considering the vertical dimension of wireless networks, aerial ISs can provide additional design DoFs for coverage enhancement. To cope with the fast time-varying channels in high-mobility scenarios, ISs can be deployed along roadside or on/in vehicles to mitigate the unfavorable effects of fast fading. For more practical use of these IS deployment strategies, however, various issues such as cross-band interference, hardware imperfections and costly channel estimation should be further investigated.
@@ II-C Modulation
From a network functionality and information transmission perspective, we can classify ISs into two types depending on their roles in the network: relay-type IS and transmitter-type IS [200]. Specifically, an individual IS can function as either a relay or a transmitter, depending on its functionality and the specific use cases. Relay-type IS, as depicted in Fig. 12 and discussed in the previous subsections, is strategically positioned between the source (e.g., BS) and the destination (e.g., users) to enhance communication performance between them, especially when the direct link is considerably weak or obstructed. On the other hand, transmitter-type IS is an integral part of the transmitter, participating in its signaling and modulation processes, as illustrated in Fig. 12. A key requirement is the capability for each IS element’s phase/amplitude change at the symbol level, which poses a critical challenge in hardware design and implementation. However, with typical symbol periods of current wireless systems like 5G/WiFi being in the order of microseconds, the switching speed of the IS element’s phase/amplitude can match the symbol transmission rate [201, 202]. A practical approach for realizing transmitter-type IS is by connecting IS with its aided transmitter via a reliable wired link, thus enabling real-time control of IS reflection on a per-symbol basis. In this subsection, we focus on the transmitter-type IS and provide an overview of two types of transmitter-type IS-based modulation: amplitude/phase modulation (APM) and index modulation (IM).
@@ II-C1 IS for Amplitude/Phase Modulation
On the other hand, the principle of transmitter-type IS is also similarly applied in backscatter communication, where a backscatter device (BD) transmits messages to a backscatter receiver (BR) by reflecting an incident signal from an external RF emitter. The adaptability of IS allows for its effective deployment in backscatter communication networks. Specifically, the fundamental principle of IS-aided backscatter communication was introduced in [212], with various types of modulation considered. The authors of [213] further analyzed cooperative IS and backscatter communication performance with different amplitude/phase modulations. The BER performance of IS-aided backscatter communication networks was examined in [214], demonstrating the effectiveness of ISs in improving the channel conditions of both source-reader and source-tag links. The closed-form achievable rate expressions for IS-aided backscatter communication networks were provided in [215]. Further investigation was made in [216], where the achievable rate of IS-aided backscatter communication networks adopting $M$-PSK was maximized by jointly optimizing the phase shifts of IS. Unlike the aforementioned literature where an IS is used solely for link enhancement, [203] explored the function of an IS as an AP with an unmodulated carrier, showing that ISs can also perform backscatter communication. Building on this result, [217] proposed dividing an IS into two parts, one for transmitting the primary data signal and the other for transmitting the backscatter signal.
@@ II-C2 IS for Index Modulation
Summary: In this subsection, we have mainly reviewed the works on transmitter-type IRS, which serves a unique role in network functionality. A transmitter-type IS can simplify hardware architecture and enhance signal quality by participating in signaling and modulation processes, which are generally based on two types of modulation: APM and IM. Specifically, ISs can implement various modulation schemes (e.g., PSK, QAM, and FSK) and act as virtual MIMO terminals, thus enabling transmit diversity. Through IM techniques, ISs can selectively activate/deactivate reflecting elements to convey additional information. Despite these promising recent advances in IS-based modulation designs, critical challenges remain to be tackled for fully unlocking their potential for future wireless networks. As the changing speed of each IS element’s phase/amplitude at the symbol level poses stringent requirements, the resulted low data rate of reflection modulation is another practical issue to address in future work.
@@ II-D Sensing and ISAC
ISAC has been regarded as a promising technique for the 6G wireless network, where the sensing function and the communication function can be carried out simultaneously by sharing the same spectrum and hardware facilities [227, 228, 229]. However, ISAC systems still face practical challenges. For example, sensing coverage is significantly smaller than communication coverage due to severe path loss in the BS-IS-target-IS-BS cascaded echo link. Although the sensing coverage can be guaranteed by deploying more BSs, this method incurs high hardware cost and energy consumption. On the other hand, hardware sharing of communication and sensing transceivers increases the design and implementation complexity, while spectrum sharing of communication and sensing signals results in interference and performance trade-off. Fortunately, the use of IS can achieve higher spectral/energy efficiency for wireless communication[230, 231], as well as higher accuracy for wireless sensing, thus promising to enable ISAC for 6G [232, 233, 234]. Specifically, the IS can establish LoS channels to enhance both sensing/communication coverage, provide flexible passive beamforming to balance their performance trade-off, and reduce hardware cost and energy consumption compared to dense deployment of BSs [235, 236]. Also, equipping ISs with sensing capabilities enables an easier deployment of the ISs [237] and energy harvesting capabilities [238].